# Conceptualization, Data Curation, Formal Analysis, Methodology, Software, Writing – original draft

### v1

@@ Abstract
Flight delay prediction has become a key focus in air traffic management, as delays highlight inefficiencies that impact overall network performance. This paper presents a lightweight large language model-based multimodal flight delay prediction, formulated from the perspective of air traffic controllers monitoring aircraft delay after entering the terminal area. The approach integrates trajectory representations with textual aeronautical information, including flight information, weather reports, and aerodrome notices, by adapting trajectory data into the language modality to capture airspace conditions. Experimental results show that the model consistently achieves sub-minute prediction error by effectively leveraging contextual information related to the sources of delay. The framework demonstrates that linguistic understanding, when combined with cross-modality adaptation of trajectory information, enhances delay prediction. Moreover, the approach shows practicality and scalability for real-world operations, supporting real-time updates that refine predictions upon receiving new operational information.
@@ 1 Introduction
The volume of air traffic grows in parallel with economic development. This growth can result in traffic density that exceeds the capacity of current air traffic management (ATM) systems, compromising both safety and operational efficiency [1, 2]. Such inefficiencies lead to flight delays that increase airline operating costs, burden airport coordination, disrupt passenger itineraries, and create network-wide ripple effects. The Federal Aviation Administration (FAA) defines delay as the excess time duration incurred during flight operations. These delays are divided into five operational phases, namely gate delay, taxi-out delay, en-route delay, terminal delay, and taxi-in delay [3]. Delays can also be classified by cause, including carrier delays, weather delays, security delays, and airspace-related delays [4]. The delays across operational phases and causes originate from different parts of the ATM system. Understanding relevant factors is essential for accurate, phase-specific delay prediction. This study focuses on delays after the terminal phase, which are of particular interest to air traffic controllers (ATCs), as monitoring and mitigating delays within this phase is their direct responsibility.
@@
In ATM operations, the ATCs monitor trajectory data and textual aeronautical information, which belong to different modalities. Although these data are comprehensively human-interpretable, managing, inferring, and making decisions based on large volumes of such multi-modal information can still be overwhelming for ATCs. The growing public availability of large-scale aviation data has enabled the development of data-driven models to assist ATCs in these tasks, enhancing autonomy while preserving effective human–machine interaction. Textual information, such as flight schedules, weather reports, weather forecasts, and temporary operational notices, provides essential context for aviation operations and is nowadays accessible via many public databases. Aircraft trajectory data, which describes aircraft movement and operational behavior within controlled airspace, can be extracted from recorded Automatic Dependent Surveillance–Broadcast (ADS-B) databases. Integrating and understanding these contexts plays an important role in improving airspace-related delay prediction.
@@
Since the operational context comprises information in different modalities, each requires distinct handling techniques, while their combined use requires effective data integration. For textual data, large language models (LLMs) have recently gained recognition as powerful models capable of understanding natural language. Pretrained on large-scale language corpora, LLMs can often interpret aviation-related text, such as flight descriptions and aeronautical information. On the other hand, trajectories are complex, high-dimensional time series that can be represented in a more compact form, which helps downstream models perform better even without further fine-tuning. However, trajectory data is not the type of input recent LLMs are trained to handle directly; therefore, integrating such data with LLMs requires additional adaptation techniques. This paper’s proposed method incorporates two pre-trained models, an LLM and a trajectory representation model, using a simple yet effective adaptation technique to integrate their embeddings. We formulate flight delay prediction as a multimodal regression problem, where both textual and trajectory data serve as inputs for prediction. The contributions of this paper are as follows:
@@
  • We propose a language-trajectory multimodal framework for predicting delays beyond the terminal phase, incorporating comprehensive operational context through the integration of textual aeronautical information and aircraft trajectories, leveraging an LLM and a trajectory representation model for their respective modalities.
@@
  • We demonstrate that incorporating a wide range of aeronautical information into the operational context leads to more accurate delay predictions, as the model can account for contextual factors that describe the situation in the airspace, which in turn extend the time an aircraft spends within the area.
@@ 2.1 Delay Prediction
Flight delay prediction has long been an area of interest in the ATM community. Early studies employed tabular-based approaches, in which flight details (e.g., carrier, route, schedule, delay history) and contemporaneous weather data (e.g., visibility, wind, temperature) were extracted into structured tables and used as fixed-dimensional numerical or categorical features for machine learning models. Delay prediction can be formulated as a classification problem, where the goal is to determine whether a delay will occur. For example, prior studies have employed logistic regression [5] and a modified Random Forest model [6]. Meanwhile, [7, 8, 9, 10] have compared various non-neural machine learning models for delay classification. Integrating traffic count features, [11] predicts departure delays using linear regression, support vector machines (SVM), Multilayer Perceptron (MLP), and LightGBM. The delay prediction can also be formulated as a regression problem that predicts delay duration in minutes. For example, SVM [12] and an MLP using only flight information [13] have been applied to predict delay duration. [14] compares several models for predicting delay distributions from flight features; subsequently, [15, 16, 9] compare various models for delay regression.
@@
Many works have incorporated historical data for delay classification in sequence. For instance, [17, 18] both employ long short-term memory (LSTM) models using flight information and weather data to classify next-day delay. However, the day-level granularity in both limits their real-time applicability. Complex network models, such as the LSTM-integrated random forest model [19] and the Convolutional Neural Network–Long Short-Term Memory (CNN-LSTM) [20], were employed to predict delays based on weather, delay propagation, and airport congestion, as measured by traffic count. [21] models multi-flight delay propagation using attention-based architectures that capture flight chain dependencies across sequential operations of the same aircraft, accounting for temporal coupling and propagation effects in aircraft rotations. [4] employs Gaussian Process Regression to predict flight delays using only the prior delays and flight dates, with the model learning the mean and covariance from historical delay patterns.
@@
Several recent studies have adopted graph-based approaches for delay regression. [22] employs a GCN-CNN architecture, where a Graph Convolutional Network (GCN) captures spatial dependencies in a time-evolving network of airports, and a CNN models daily and weekly delay patterns. Similarly, [23] proposes a hybrid framework that combines diffusion GCNs with residual Gated Recurrent Units (GRUs), incorporating federated learning across airport networks for accurate, privacy-preserving forecasting using historical delay data. [24] represents air traffic as a directed multigraph for predicting delays caused by holding maneuvers. [25] employs an attention-based LSTM model combining aircraft trajectory with flight and weather data to classify delays as early, on-time, or late.
@@ 2.3 Time Series and Multimodal Models
2.4 Large Language Model for Aviation Applications
@@ 3.1 Problem Formulation
$D_{pre}$ $=D_{gate}+D_{taxi-out}+\left(T_{entry}^{act}-T_{entry}^{exp}-D_{gate}-D_{taxi-out}\right)$ (2)
@@
$D_{post}$ $=\left(T_{land}^{act}-T_{land}^{exp}\right)-\left(T_{entry}^{act}-T_{entry}^{exp}\right)+\left(T_{arr}^{act}-T_{arr}^{sch}\right)-\left(T_{land}^{act}-T_{land}^{exp}\right)$ (5)
@@
$D_{total}$ $=\left(T_{entry}^{act}-T_{entry}^{exp}\right)+\left(T_{arr}^{act}-T_{arr}^{sch}\right)-\left(T_{entry}^{act}-T_{entry}^{exp}\right)$ (8)
@@
A scenario, $s_{i,t}$, comprises data from multiple modalities. The textual prompt $P^{fp}_{t}$ encodes general flight information, $P^{wx}_{t}$ captures weather conditions by combining METAR and TAF reports, and $P^{no}_{t}$ contains NOTAMs information. This paper introduces a trajectory-based characterization of airspace conditions, enabling the model to interpret congestion more effectively than the traffic count methods used in [11, 19, 20]. The trajectory data is represented as a time-series modality, where each timestep contains the $x,y,z$ coordinates in the ENU (East-North-Up) frame. To describe the airspace, we use three distinct groups of trajectories. The focusing trajectory $X^{f}_{i,t}$, the active trajectories $X^{a}_{i,t}$ those currently airborne at time $t$, and the completed prior trajectories $X^{p}_{i,t}$, which may provide additional context or relevant information. The prediction model is formulated as $\hat{y}_{i}=f_{\theta}(s_{i,t})$, where $\hat{y}_{i}\equiv\Delta\hat{t}_{post,i}$, and the two notations are used interchangeably. The function $f_{\theta}$ represents the overall model, composed of several components. The text part includes a tokenizer $T(\cdot)$, followed by an embedding table $\text{Emb}[\cdot]$. The trajectory part consists of the ATSCC representation network, denoted as $f_{\text{atscc}}(\cdot)$ [56], followed by a trainable cross-modality adaptation network $f_{\text{xa}}(\cdot)$. The combined embeddings are then passed to a frozen LLM backbone $f_{\text{llm}}(\cdot)$. Finally, a trainable prediction head $f_{h}(\cdot)$ that predicts $\hat{y}_{i}$. The training of the model $f_{\theta}(\cdot)$ is formulated as a regression problem, minimizing the error between the predicted post-terminal duration $\hat{y}_{i}$ and the ground-truth $y_{i}$.
@@ 3.2.1 General Flight Information Processing
We first construct a flight schedule database using data on arrival flights at Incheon International Airport for the year 2022, queried from Airportal [75]. The retrieved records consist of structured information, including scheduled arrival times, airline names, flight identifiers, and airport details such as departure and destination codes, as well as airport names. We matched the airports with their coordinates, including latitude, longitude, and altitude, and calculated the great-circle distance between the origin and destination. The flight haul type, categorized as short-haul, medium-haul, or long-haul, is determined based on this distance according to the classification provided in Wragg’s aviation dictionary [76]. The date and day of the week of arrival are recorded in Korea Standard Time. Certain fields are synthesized to complete the dataset. Aircraft type and registration are inferred based on the frequently used aircraft models on each specific route. The wake turbulence category is classified based on the aircraft type in accordance with ICAO’s Procedures for Air Navigation Services - Air Traffic Management Doc 4444 [77]. The actual airspace entry time and the current time stamp ($t$) are included here to provide additional context. Each flight $i$ is associated with general flight information, as detailed in Table 2, which is later used to construct the prompt $P^{fp}_{i,t}$ using the format in Figure 9.
@@
Flight Information Description Data Type
@@
haul Flight haul type (short, medium, long, local) string
@@
dep_lat Departure airport latitude float64
dep_lon Departure airport longitude float64
dep_altitude Departure airport altitude (ft) float64
@@
dest_lat Destination airport latitude float64
dest_lon Destination airport longitude float64
dest_altitude Destination airport altitude (ft) float64
distance Great-circle distance between origin and destination (km) float64
actual_entry_time Time of first ADS-B appearance or airspace entry datetime64
sched_time_utc Scheduled arrival time (UTC) datetime64
date Scheduled Flight Date of arrival (KST) datetime64
@@ 3.2.3 Trajectory Data Preprocessing
  • Prior trajectories: Using the earliest timestamp among all active trajectories as a reference point, we collect all completed prior trajectories that were active at that time. Although these flights have completed their operations, they may still hold semantic importance. They capture patterns that may influence the paths of focusing and active flights, such as when aircraft are instructed to follow earlier ones. They also indirectly visualize ground traffic activity. The prior trajectories are nan-padded and represented as $X^{p}_{i,t}\in\mathbb{R}^{N^{p}\times\mathcal{T}^{p}_{\max}\times 3}$, where $N^{p}$ denotes the number of prior trajectories at time $t$, and $\mathcal{T}^{p}_{\max}$ is the longest trajectory length among them.
Therefore, the trajectories $X^{f}_{i,t}$, $X^{a}_{i,t}$, and $X^{p}_{i,t}$ represent the query at time $t$ for flight $i$ under scenario $s_{i,t}$. Intuitively, these trajectories enable the model to achieve a comprehensive semantic understanding of the airspace, akin to the mental image used by human ATCs, without requiring explicit procedures such as trajectory classification or waypoint capacity estimation. This allows the model to directly perform the target task, estimating the post-terminal duration.
@@ 3.2.4 Label Preparation
The corresponding label $y_{i}$ for the regression task associated with $s_{i,t}$ was constructed as part of the supervised training setup. For flight $i$, the actual post-terminal duration, denoted as $\Delta t_{post,i}$, is calculated using the airspace entry time $T_{entry,i}^{act}$, derived from ADS-B data [80], taken as the moment the aircraft enters Incheon Airport’s TMA, and the actual arrival time $T_{arr,i}^{act}$, obtained from the official records in hour–minute format in Airportal [75]:

### v2

# Conceptualization, Data Curation, Formal Analysis, Methodology, Software, Writing - original draft

@@ Abstract
Flight delay prediction has become a key focus in air traffic management, as delays highlight inefficiencies that impact overall network performance. This paper presents a lightweight large language model-based multimodal flight delay prediction, formulated from the perspective of air traffic controllers monitoring aircraft delay after entering the terminal area. The approach integrates trajectory representations with textual aeronautical information, including flight information, weather reports, and aerodrome notices, by adapting trajectory data into the language modality to capture airspace conditions. The experiments show that the model consistently achieves sub-minute prediction error by effectively leveraging contextual information related to the sources of delay, fulfilling the operational standard for minute-level precision. The framework demonstrates that linguistic understanding, when combined with cross-modality adaptation of trajectory data, enhances delay prediction. Moreover, the approach shows practicality and potential scalability for real-world operations, supporting real-time updates that refine predictions upon receiving new operational information.
@@ 1 Introduction
The volume of air traffic grows in parallel with economic development. This growth can result in traffic density that exceeds the capacity of current air traffic management (ATM) systems, compromising both safety and operational efficiency [1, 2]. Such inefficiencies lead to flight delays that increase airline operating costs, burden airport coordination, disrupt passenger itineraries, and create network-wide ripple effects. The Federal Aviation Administration (FAA) defines delay as the excess time duration incurred during flight operations. These delays are categorized into five operational phases, namely gate delay, taxi-out delay, en-route delay, terminal delay, and taxi-in delay [3]. Delays can also be classified by cause, including carrier delays, weather delays, security delays, and airspace-related delays [4]. The delays across operational phases and causes originate from different parts of the ATM system. Understanding relevant factors is essential for accurate, phase-specific delay prediction. This study focuses on delays after the terminal phase, which are of particular interest to air traffic controllers (ATCs), as monitoring and mitigating delays within this phase is their direct responsibility.
@@
In ATM operations, the ATCs monitor trajectory data and textual aeronautical information, which belong to different modalities. Although these data are comprehensible and human-interpretable, managing, inferring, and making decisions from large volumes of such multimodal information can still be overwhelming for ATCs. The growing public availability of large-scale aviation data has enabled the development of data-driven models to assist ATCs in these tasks, enhancing autonomy while preserving effective human-machine interaction. Textual information, such as flight schedules, weather reports, and temporary operational notices, provides essential context for aviation operations and is nowadays accessible via many public databases. Aircraft trajectory data, which describes aircraft movement and operational behavior within controlled airspace, can be extracted from recorded Automatic Dependent Surveillance-Broadcast (ADS-B) databases. Integrating and understanding these contexts is important for capturing potential delay-related factors in the airspace.
@@
Since the operational context comprises information in various modalities, each requires distinct handling techniques, while their combined use necessitates effective data integration. For textual data, large language models (LLMs) have recently gained recognition as powerful models capable of understanding natural language. Pretrained on large-scale language corpora, LLMs can often interpret aviation-related text, such as flight descriptions and aeronautical information. On the other hand, trajectories are complex, high-dimensional time series that can be represented more compactly, helping downstream models perform better even without further fine-tuning. However, trajectory data is not the type of input recent LLMs are trained to handle directly; therefore, integrating such data with LLMs requires additional adaptation techniques. This paper’s proposed method incorporates two pre-trained models, an LLM and a trajectory representation model, using a simple yet effective adaptation technique to integrate their embeddings. We formulate flight delay prediction as a multimodal regression problem, where both textual and trajectory data serve as context. The contributions of this paper are as follows:
@@
  • We propose a language-trajectory multimodal framework for predicting delays beyond the terminal phase, incorporating comprehensive operational context by integrating textual aeronautical information and aircraft trajectories, leveraging an LLM and a trajectory representation model for their respective modalities.
@@
  • We demonstrate that incorporating a wide range of aeronautical information into the operational context leads to more accurate delay predictions, as the model can account for contextual factors that describe the airspace, thereby extending the time an aircraft spends within the area.
@@ 2.1 Delay Prediction
Flight delay prediction has long been an area of interest in the ATM community. Early studies employed tabular-based approaches, in which flight details (e.g., carrier, route, schedule, delay history) and contemporaneous weather data (e.g., visibility, wind, temperature) were extracted into structured tables and used as fixed-dimensional numerical or categorical features for machine learning models. Delay prediction can be formulated as a classification problem, where the goal is to determine whether a delay will occur. For example, prior studies have employed logistic regression [5] and a modified Random Forest model [6]. Meanwhile, several studies [7, 8, 9, 10] have compared various non-neural machine learning models for delay classification. Integrating traffic count features, a model presented in [11] predicts departure delays using linear regression, support vector machines (SVM), Multilayer Perceptron (MLP), and LightGBM. The delay prediction can also be formulated as a regression problem that predicts delay duration in minutes. For example, SVM [12] and an MLP using only flight information [13] have been applied to predict delay duration. The work in [14] compares several models for predicting delay distributions based on flight features. Subsequently, studies [15, 16, 9] compare various models for delay regression.
@@
Many works have incorporated historical data for delay classification in sequence. For instance, the methods proposed in [17, 18] both employ long short-term memory (LSTM) models, utilizing flight information and weather data to classify next-day delays. However, the day-level granularity in both limits their real-time applicability. Complex network models, such as the LSTM-integrated random forest model [19] and the Convolutional Neural Network-Long Short-Term Memory (CNN-LSTM) [20], were employed to predict delays based on weather, delay propagation, and airport congestion, as measured by traffic count. Qu et al. [21] model multi-flight delay propagation using attention-based architectures that capture flight chain dependencies across sequential operations of the same aircraft, accounting for temporal coupling and propagation effects in aircraft rotations. Khanal et al. [4] employ Gaussian Process Regression to predict flight delays using only the prior delays and flight dates, with the model learning the mean and covariance from historical delay patterns.
@@
Several recent studies have adopted graph-based approaches for delay regression. Cai et al. [22] employ a GCN-CNN architecture, where a Graph Convolutional Network (GCN) captures spatial dependencies in a time-evolving network of airports, and a CNN models daily and weekly delay patterns. Similarly, a hybrid framework, proposed in [23], combines diffusion GCNs with residual Gated Recurrent Units (GRUs) and incorporates federated learning across airport networks to achieve accurate, privacy-preserving forecasting using historical delay data. Franco et al. [24] model air traffic as an airport multigraph to predict delays caused by holding maneuvers. A recent trajectory-based approach presented in [25] employs an attention-based LSTM model that combines aircraft trajectory data with flight and weather data to classify delays as early, on-time, or late.
@@ 2.3 Time Series and Multimodal Models
2.4 Large Language Models for Aviation Domain
@@ 3.1 Problem Formulation
$D_{pre}$ $=D_{gate}+D_{taxi-out}$
$+\left(T_{entry}^{act}-T_{entry}^{exp}-D_{gate}-D_{taxi-out}\right)$ (2)
@@
$D_{post}$ $=\left(T_{land}^{act}-T_{land}^{exp}\right)-\left(T_{entry}^{act}-T_{entry}^{exp}\right)$
$+\left(T_{arr}^{act}-T_{arr}^{sch}\right)-\left(T_{land}^{act}-T_{land}^{exp}\right)$ (5)
@@
$D_{total}$ $=\left(T_{entry}^{act}-T_{entry}^{exp}\right)+\left(T_{arr}^{act}-T_{arr}^{sch}\right)$
$-\left(T_{entry}^{act}-T_{entry}^{exp}\right)$ (8)
@@
A scenario, $s_{i,t}$, comprises data from multiple modalities. The textual prompt $P^{fp}_{t}$ encodes general flight information, $P^{wx}_{t}$ captures weather conditions by combining METAR and TAF reports, and $P^{no}_{t}$ contains NOTAMs information. This paper introduces a trajectory-based characterization of airspace conditions, enabling the model to interpret congestion more effectively than traffic-count methods used in [11, 19, 20]. The trajectory data is represented as a time series, where each timestep contains the $x,y,z$ coordinates in the ENU (East-North-Up) frame. To describe the airspace, we use three distinct trajectory groups. The focusing trajectory $X^{f}_{i,t}$, the active trajectories $X^{a}_{i,t}$ those currently airborne at time $t$, and the completed prior trajectories $X^{p}_{i,t}$, which may provide additional context or relevant information. The prediction model is formulated as $\hat{y}_{i}=f_{\theta}(s_{i,t})$, where $\hat{y}_{i}\equiv\Delta\hat{t}_{post,i}$, and the two notations are used interchangeably. The function $f_{\theta}$ represents the overall model, composed of several components. The text part includes a tokenizer $T(\cdot)$, followed by an embedding table $\text{Emb}[\cdot]$. The trajectory part consists of the ATSCC representation network, denoted as $f_{\text{atscc}}(\cdot)$ [56], followed by a trainable cross-modality adaptation network $f_{\text{xa}}(\cdot)$. The combined embeddings are then passed to a frozen LLM backbone $f_{\text{llm}}(\cdot)$. Finally, a trainable prediction head $f_{h}(\cdot)$ that predicts $\hat{y}_{i}$. The training of the model $f_{\theta}(\cdot)$ is formulated as a regression problem, minimizing the error between the predicted post-terminal duration $\hat{y}_{i}$ and the ground-truth $y_{i}$.
@@ 3.2.1 General Flight Information Processing
We first construct a flight schedule database using arrival flight data for 2022 at Incheon International Airport, obtained from Airportal [75]. The retrieved records consist of structured information, including scheduled arrival times, airline names, flight identifiers, and airport details such as departure and destination codes, as well as airport names. We matched the airports with their coordinates, including latitude, longitude, and altitude, and calculated the great-circle distance between the origin and destination. The flight haul type, categorized as short-haul, medium-haul, or long-haul, is determined based on this distance according to the classification provided in Wragg’s aviation dictionary [76]. The date and dayarrival date and day of the week in Korea Standard Time. Certain fields are synthesized to complete the dataset. Aircraft type and registration are inferred based on the frequently used aircraft models on each specific route. The wake turbulence category is classified based on the aircraft type in accordance with ICAO’s Procedures for Air Navigation Services - Air Traffic Management Doc 4444 [77]. The actual airspace entry time and the current time stamp ($t$) are included here to provide additional context. Each flight $i$ is associated with general flight information, as detailed in Table 2, which is later used to construct the prompt $P^{fp}_{i,t}$ using the format in Figure 9.
@@
Flight Information Description Type
@@
haul Flight haul type string
@@
dep_lat Departure airport latitude float
dep_lon Departure airport longitude float
dep_altitude Departure airport altitude (ft) float
@@
dest_lat Destination airport latitude float
dest_lon Destination airport longitude float
dest_altitude Destination airport altitude (ft) float
distance Great-circle distance (km) float64
actual_entry_time Time of first ADS-B appearance (UTC) datetime
sched_time_utc Scheduled arrival time (UTC) datetime
date Scheduled Flight Date of arrival (KST) datetime
@@ 3.2.3 Trajectory Data Preprocessing
  • Prior trajectories: Using the earliest timestamp among all active trajectories as a reference point, we collect all completed prior trajectories that were active at that time. Although these flights have completed their operations, they may still hold semantic importance. They capture patterns that may influence the paths of focusing and active flights, such as when aircraft are instructed to follow earlier ones. They also indirectly visualize ground traffic. The prior trajectories are nan-padded and represented as $X^{p}_{i,t}\in\mathbb{R}^{N^{p}\times\mathcal{T}^{p}_{\max}\times 3}$, where $N^{p}$ denotes the number of prior trajectories at time $t$, and $\mathcal{T}^{p}_{\max}$ is the longest trajectory length among them.
@@
The trajectories $X^{f}_{i,t}$, $X^{a}_{i,t}$, and $X^{p}_{i,t}$ represent the query at time $t$ for flight $i$ under scenario $s_{i,t}$. Intuitively, these trajectories enable the model to achieve a comprehensive semantic understanding of the airspace, akin to the mental image used by human ATCs, without requiring explicit procedures such as trajectory classification or waypoint capacity estimation. This allows the model to directly perform the target task, estimating the post-terminal duration.
Figure 3: Dataset Statistics
@@ 3.2.4 Label Preparation
The corresponding label $y_{i}$ for the regression task associated with $s_{i,t}$ was constructed as part of the supervised training setup. For flight $i$, the actual post-terminal duration, denoted as $\Delta t_{post,i}$, is calculated using the airspace entry time $T_{entry,i}^{act}$, derived from ADS-B data [80], taken as the moment the aircraft enters Incheon Airport’s TMA, and the actual arrival time $T_{arr,i}^{act}$, obtained from the official records in hour-minute format in Airportal [75]: